# ResQ: Mixed-Precision Quantization of Large Language Models with Low-Rank Residuals

### v1

@@ 4.3 Inference computation with optimized projections
Once the projection matrices are obtained, the operation defined in Equation 4 requires multiplying the weights and activations with ${{\bm{U}}}$. Weights can be projected and quantized offline. The projection operation on an activation can be merged to the weight of a previous linear layer. Based on the architecture of decoder based LLMs, we introduce four different kinds of projections (Figure 3) : ${{\bm{U}}_{A}\in{\mathbb{R}}^{d_{h}\times d_{h}}}$, ${{\bm{U}}_{B},{\bm{U}}_{C}\in{\mathbb{R}}^{d_{\text{head}}\times d_{\text{head}}}}$, ${{\bm{U}}_{D}\in{\mathbb{R}}^{d_{\text{FFN}}\times d_{\text{FFN}}}}$ where $d_{h}$ is hidden dimension of LLM, $d_{\text{head}}$ is the attention head dimension and $d_{\text{FFN}}$ the hidden dimension of feedforward network.
Projections at block boundaries  Input activations to attention and feed forward blocks are projected via ${\bm{U}}_{A}$. Projection is handled by post multiplying the weight matrix of final linear layer in each block (o_proj in attention and down_proj in feed forward) by ${\bm{U}}_{A}$. Thus, projections of activations is handled at no additional inference cost. To maintain numerical invariance, the first linear layer of each block (q|k|v_proj in attention and up_proj|gate_proj in feed forward) is pre multiplied with ${{\bm{U}}_{A}^{\top}}$ (Figure 3 a). Similarly, the weights of the embedding layer and the final head are modified to manage projection of the residual stream.
Family Method Training Llama-2-7b Llama-2-13b
@@
16-bit 5.5 64.1 41.1 4.9 66.5 52.7
[…]
GPTQ 9600.0 38.9 23.8 3120.0 33.8 24.8
SmoothQuant+ 15.4 46.1 24.2 11.2 51.0 26.6
QUIK 7.5 57.0 27.8 6.8 60.2 36.7
QuaRot 6.1 60.7 32.3 5.4 63.8 47.7
SpinQuant 6.0 61.0 34.8 5.2 64.8 47.8
Llama-2 ResQ 5.8 62.0 37.7 5.1 65.2 50.1
Family Method Training Llama-3-8b Llama-3-70b
16-bit 6.1 67.1 63.1 2.9 73.1 75.9
@@
Llama-3.2 ResQ 12.4 50.1 29.4 8.8 59.0 49.8
Table 1: Comparison of perplexity score on Wikitext, average zero-shot common sense reasoning accuracy and average zero-shot MMLU accuracy at W/A/KV = 4/4/4. Results of all techniques were obtained using their official codebase. Our work ResQ and QUIK Ashkboos et al. (2023) keep $\nicefrac{{1}}{{8}}$th channels in 8-bit. All techniques except RTN use GPTQ Frantar et al. (2022) for weight quantization. $\uparrow$ higher is better, $\downarrow$: lower is better. Full results in App. A.3
@@
${{\bm{q}}_{\text{proj}}{\bm{K}}^{\top}_{\text{proj}}=({\bm{q}}{\bm{U}}_{C})({\bm{U}}_{C}^{\top}{\bm{K}}^{T})={\bm{q}}{\bm{K}}^{\top}}.$ (8)
Where ${\bm{q}}$ and ${\bm{K}}$ are query and key after rotary embedding (RoPE) respectively. Unfortunately, ${{\bm{U}}_{C}}$ cannot be merged into the previous linear layer due to presence of RoPE and the projection is applied on the fly. To make the on the fly projection efficient, we apply uniform precision quantization to ${\bm{U}}_{C}$ and corresponding input activations.
Projections within the feedforward block  ${{\bm{U}}_{D}}$ ensure improved quantization of activation within feed forward block (Figure 3c). ${\bm{U}}^{\top}_{D}$ is pre-multiplied with weights of down_proj while due to the presence of activation function within the block, ${{\bm{U}}_{D}}$ cannot be merged to weights of preceeding linear layer and is left on the fly as well. ${\bm{U}}_{D}$ is applied to the hidden dimension of the feed forward network ($d_{\text{FFN}}$) which is 3 to 4 times greater than the hidden dimension of LLM in most language models. In this scenario, matrix multiplication with ${\bm{U}}_{D}$ is extremely expensive in computation and storage. To minimize the overhead, we choose ${\bm{U}}_{D}$ to be a hadamard matrix to leverage fast and efficient hadamard transform kernel. And, we choose weights and activations for down_proj layer to be uniformly quantized to low precision.
@@
Table 2: Comparison of performance of quantization approaches on generative tasks at precisions of W/A/KV = 4/4/4 bits. Our work ResQ and QUIK Ashkboos et al. (2023) keep $\nicefrac{{1}}{{8}}$th of channels in 8-bit.
@@ 5.1 Setup
Models, tasks, datasets and baselines  We conduct experiments on Llama 2 Touvron et al. (2023), Llama 3 Meta (2024a) and very recent Llama 3.2 Meta (2024b) models. We benchmark our approach against GPTQ Frantar et al. (2022), QuaRot Ashkboos et al. (2024b), QUIK Ashkboos et al. (2023), SpinQuant Liu et al. (2024a) and SmoothQuant+, a stronger baseline created by combining SmoothQuant Xiao et al. (2023) with GPTQ following Sharify et al. (2024). We evaluate the quantization approaches on a range of tasks which measure the language modeling ability: perplexity on Wikitext Merity et al. (2016), common sense reasoning ability: average 0-shot accuracy on Arc-c/e Clark et al. (2018), BoolQ Clark et al. (2019), HellaSwag Zellers et al. (2019), Openbook QA Mihaylov et al. (2018), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), WinoGrande Sakaguchi et al. (2021), language understanding: 0-shot accuracy on MMLU Hendrycks et al. (2021), mathematical understanding: 5-shot GSM8K Cobbe et al. (2021), dialogue summarization: samsum Gliwa et al. (2019) and qmsum Zhong et al. (2021) from LongBench Bai et al. (2024), and code completion: repobench-p Liu et al. (2023b) from LongBench.
Implementation details  We implement ResQ using the HuggingFace Transformers library Wolf et al. (2020) with PyTorch Paszke et al. (2019). We share a single ${{\bm{U}}_{A}}$ across all layers, while ${{\bm{U}}_{B}}$, ${{\bm{U}}_{C}}$ and ${{\bm{U}}_{D}}$ are generated per layer. Following SpinQuant Liu et al. (2024a), we use per-token asymmetric quantization for activations, per-channel symmetric quantization for weights, and per-head asymmetric quantization for the KV cache. We fuse the projection matrices ${{\bm{U}}_{A}},{{\bm{U}}_{B}},{{\bm{U}}_{D}}$ into weights and apply GPTQ Frantar et al. (2022) for weight quantization. To efficiently implement on-the-fly projections, ${{\bm{U}}_{D}}$ is a Hadamard matrix and ${\bm{U}}_{C}$ and its corresponding activations are quantized to 8-bit. The entire process, including obtaining projections and quantization, runs on a single NVIDIA A100 GPU; for Llama-3-8B, it takes 35 minutes. Additional details are in Appendix A.4.
W/A/KV Method Llama-3-8b Llama-3.2-3b
@@
Table 3: Wikitext perplexity comparison of ResQ and baseline which keeps channels with high $l_{\infty}$ norm in high precision and uses rotation to reduce quantization error within high precision and low precision groups.
@@ 5.2 Main Results
Language modeling, understanding, and reasoning tasks  We evaluate ResQ on tasks that test language modelling ability (Perplexity on Wikitext), common sense reasoning ability (average 0-shot accuracy on the eight tasks listed in section 5.1) and language understanding (average 0-shot accuracy on MMLU). The results are presented in Table 1. We see that ResQ reduces the gap to 16-bit performance and outperforms the quantization baselines across all tasks on all models. Particularly, ResQ outperforms SpinQuant by achieving 2-33% lower Wikitext perplexity, 0.3-5.4% better average zero-shot accuracy and a 1-14.5% better accuracy on MMLU benchmark across the models tested without any additional training. Compared with QUIK, another mixed precision quantization approach, ResQ achieves 22-50% better perplexity, 5-12.3% better average zero shot accuracy and 4.3-24.5% better MMLU accuracy over all models.
Generative tasks  We also test ResQ on tasks that require auto-regressive token generation including the GSM8K mathematical understanding benchmark, dialogue summarization benchmarks (qmsum and samsum) and code completion benchmark (repobench-p) (in Table 2). The goal of choosing these tasks is to evaluate the generation ability on a wide variety of domains. On the challenging GSM8K benchmark where QUIK fails to produce meaningful results, ResQ outperforms SpinQuant by 3.8% and 5.5% on the 8b and 3b parameter model respectively, closing the gap to 16-bit baseline. On Long Bench evaluation tasks, ResQ demonstrates competitive performance and outperforms SpinQuant without any additional training.
Comparison against outliers with rotation baseline  A stronger baseline can be created combining existing quantization approaches. Like QUIK, one can find channels which consistently contain outliers and keep them in 8-bit while keep the remaining channels in low precision. And, the quantization of high/low precision groups can be improved using random rotations introduced in QuaRot. Compared with such a baseline which keeps channels with high $l_{\infty}$ norm in 8-bit, ResQ’s unique approach involves keeping coefficients along basis vectors with high eigenvalues in 8-bit. We see in Table 3 that ResQ consistently outperforms such a strong baseline across various precisions of W/A/KV highlighting ResQ’s PCA driven theoretically optimal approach of choosing high precision components.
@@
Table 4: Impact of different projections in ResQ. Evaluated by removing individual components and observing Wikitext perplexity.
@@ 5.3 Hardware performance
We implement the mixed-precision quantization using CUDA 11.8 and PyTorch. We use CUTLASS Thakkar et al. (2023) to perform int4 and int8 GEMM operations on TensorCore. We achieve about 1.25×1.25$\times$1.25 × to 2.4×2.4$\times$2.4 × speedup with ResQ over 16-bit baseline tested across several layer dimensions found in Llama models tested on NVIDIA A100 GPUs (Figure 4). We observe that larger layers achieve higher speedup with ResQ.
Figure 4: Layerwise speedup with ResQ on NVIDIA A100 over 16-bit baseline
@@ 5.4 Ablation studies
Projection bases  We evaluate the impact of different projections employed in ResQ by removing them and evaluating performance in Table 4. We see that removing ${{\bm{U}}_{D}}$ or ${{\bm{U}}_{A}}$ has a catastrophic impact on perplexity highlighting the importance of the projections. ${{\bm{U}}_{B}}$ and ${{\bm{U}}_{C}}$ which aid in quantization of KV cache have less severe impact when removed independently. But removing both of them leads to a non trivial increase in perplexity (particularly for Llama-3-8b and Llama-3.2-3b which employ grouped query attention).
@@
Calibration dataset  We change number of Wikitext calibration samples used to obtain projections and evaluate performance in Figure 5(b). We see that for Llama-3-8b, MMLU accuracy increases with increasing samples and saturates beyond 128 samples. For Llama-3.2-3b, the trend is unclear with 512 samples achieving best performance.
@@ 6 Conclusion
We introduce ResQ, a novel mixed-precision PTQ technique toward 4-bit quantization of large language models. ResQ projects weight, activation, and KV cache tensors to subspaces spanned by principal components, quantizing a low-rank ($\nicefrac{{1}}{{8}}$ of hidden dimension) high-variance subspace to 8-bit and the rest to 4-bit. ResQ achieves superior performance to both uniform- and mixed-precision quantization methods. We demonstrate the effectveness of ResQ across a variety of tasks—including language modeling, language understanding, common-sense reasoning, and language generation—using the Llama families of models. Compared to SpinQuant, the most competitive baseline, ResQ achieves up to 33% lower perplexity on the WikiText dataset without requiring any additional training and offers up to 2.4×2.4$\times$2.4 × speedup over the 16-bit baseline.
Limitations
ResQ takes a significant step towards efficiently serving LLMs in resource-constrained, on-device scenarios, potentially expanding the application space for these models. While we demonstrate our approach using the Wikitext Merity et al. (2016) dataset, we have not yet extensively explored whether other datasets might yield projection matrices that better optimize performance on downstream tasks. Additionally, we maintain a constant rank $r$ for the high-precision subspace across all layers, which may not be optimal. Identifying the ideal rank for each layer to balance accuracy and efficiency remains a promising direction for future work. Moreover, further optimization of projection matrices using a training based approach proposed in SpinQuant Liu et al. (2024a) would extend the improvements with our approach. But the resource intensive nature of such an approach has prohibited us from exploring the potential direction. Finally, although our approach aims to make LLMs more accessible and widespread, it does not address the potential risks of misuse for malicious purposes. To mitigate these risks, a strong commitment to user data protection, clear ethical guidelines, and transparency mechanisms is essential.
Acknowledgment

### v2

@@ 4.3 Inference computation with optimized projections
Once the projection matrices are obtained, the operation in Equation 4 requires multiplying the weights and activations with ${{\bm{U}}}$. Weights can be projected and quantized offline. The projection operation on an activation can be merged to the weight of a previous linear layer. Based on the architecture of decoder based LLMs, we introduce four different kinds of projections (Figure 4) : ${{\bm{U}}_{A}\in{\mathbb{R}}^{d_{h}\times d_{h}}}$, ${{\bm{U}}_{B},{\bm{U}}_{C}\in{\mathbb{R}}^{d_{\text{head}}\times d_{\text{head}}}}$, ${{\bm{U}}_{D}\in{\mathbb{R}}^{d_{\text{FFN}}\times d_{\text{FFN}}}}$ where $d_{h}$ is hidden dimension of LLM, $d_{\text{head}}$ is the attention head dimension and $d_{\text{FFN}}$ the hidden dimension of feedforward network (FFN).
Projections at block boundaries  Input activations to attention and FFN are projected via ${\bm{U}}_{A}$. Projection is handled by right-multiplying the weight matrix of final linear layer in each block (o_proj in attention and down_proj in FFN) by ${\bm{U}}_{A}$. Thus, projections of activations is handled at no additional inference cost. To maintain numerical invariance, the first linear layer of each block (q_proj|k_proj|v_proj in attention and up_proj|gate_proj in FFN) is pre multiplied with ${{\bm{U}}_{A}^{\top}}$ (Figure 4a). Similarly, the weights of the embedding layer and the final head are modified to manage projection of the residual stream.
Table 1: Comparison of perplexity score on Wikitext, average 0-shot common sense reasoning accuracy and average 0-shot MMLU accuracy at W/A/KV = 4/4/4 bits. Results of all techniques were obtained using their official codebase. Our work ResQ and QUIK (Ashkboos et al., 2024b) keep $\nicefrac{{1}}{{8}}$ channels in 8-bit. All techniques except RTN employ GPTQ (Frantar et al., 2022) for weight quantization. $\uparrow$ higher is better, $\downarrow$: lower is better. Full results in Appendix D, Tables 7 and 8.
Family Method Training Meta-Llama-3-8B Meta-Llama-3-70B
@@
16-bit baseline 6.1 67.1 63.1 2.9 73.1 75.9
@@
Llama 3.2 ResQ 12.4 50.1 29.4 8.8 59.0 49.8
Family Method Training Qwen2.5-3B Qwen2.5-72B
16-bit baseline 8.0 63.8 66.1 3.9 73.4 84.3
[…]
GPTQ 9977.8 35.1 23.2 37967.2 34.5 23.3
SmoothQuant+ 73306.7 34.8 23.9 - - -
QUIK 15.5 51.2 39.4 8.3 61.9 69.3
QuaRot 68.8 47.7 28.9 4.9 70.3 80.1
Qwen2.5 ResQ 9.0 61.1 61.2 4.6 72.0 81.5
@@
${{\bm{q}}_{\text{proj}}{\bm{K}}^{\top}_{\text{proj}}=({\bm{q}}{\bm{U}}_{C})({\bm{U}}_{C}^{\top}{\bm{K}}^{\top})={\bm{q}}{\bm{K}}^{\top}},$ (8)
where ${\bm{q}}$ and ${\bm{K}}$ are query and key after rotary embedding (RoPE), respectively. Because ${{\bm{U}}_{C}}$ cannot be merged into the previous linear layer due to presence of RoPE, the projection is explicitly computed at runtime, but made more efficient by applying uniform precision quantization to ${\bm{U}}_{C}$ and corresponding input activations.
Projections within the feedforward block  ${{\bm{U}}_{D}}$ ensure improved quantization of activation within FFNs (Figure 4c). ${\bm{U}}^{\top}_{D}$ is left-multiplied with weights of down_proj, but due to the presence of activation functions within the block, ${{\bm{U}}_{D}}$ cannot be merged to weights of preceding linear layers and is computed at runtime. ${\bm{U}}_{D}$ is applied to the hidden dimension of the FFNs ($d_{\text{FFN}}$) which is typically 3$\times$ to 4$\times$ the embedding dimension in most LLMs. In this scenario, matrix multiplication with ${\bm{U}}_{D}$ is extremely expensive in computation and storage. To minimize the overhead, we choose ${\bm{U}}_{D}$ to be a hadamard matrix to leverage fast and efficient hadamard transform kernel. And, we choose weights and activations for down_proj layer to be uniformly quantized to low precision.
Table 2: Comparison of performance of quantization approaches on generative tasks at precisions of W/A/KV = 4/4/4 bits. Our work ResQ and QUIK (Ashkboos et al., 2024b) keep $\nicefrac{{1}}{{8}}$ of channels in 8-bit.
@@ 5.1 Setup
Models, tasks, datasets and baselines  We conduct experiments on Llama 2 (Touvron et al., 2023), Llama 3 (Meta, 2024b), and the recently released Llama 3.2 (Meta, 2024a) and Qwen2.5 (Yang et al., 2024a) models. We also include multi-modal language models belonging to Qwen2 VL family (Wang et al., 2024) for our evaluations. We benchmark our approach against GPTQ (Frantar et al., 2022), QuaRot (Ashkboos et al., 2024c), QUIK (Ashkboos et al., 2024b), SpinQuant (Liu et al., 2024b) and SmoothQuant+, a stronger baseline created by combining SmoothQuant (Xiao et al., 2023) with GPTQ following Sharify et al. 2024. We evaluate the quantization approaches on a range of tasks which measure the language modeling ability: perplexity on Wikitext (Merity et al., 2016), common sense reasoning ability: average 0-shot accuracy on Arc-c/e (Clark et al., 2018), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), Openbook QA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), WinoGrande (Sakaguchi et al., 2021), language understanding: 0-shot accuracy on MMLU (Hendrycks et al., 2021), mathematical understanding: 5-shot GSM8K (Cobbe et al., 2021), dialogue summarization: samsum (Gliwa et al., 2019) and qmsum (Zhong et al., 2021) from LongBench (Bai et al., 2024), code completion: repobench-p (Liu et al., 2023b) from LongBench, and multi-modal understanding: MMMU (Yue et al., 2024).
Implementation details  We implement ResQ using the HuggingFace Transformers library (Wolf et al., 2020) with PyTorch (Paszke et al., 2019). We share a single ${{\bm{U}}_{A}}$ across all layers, while ${{\bm{U}}_{B}}$, ${{\bm{U}}_{C}}$ and ${{\bm{U}}_{D}}$ are generated per layer. Following SpinQuant (Liu et al., 2024b), we use per-token asymmetric quantization for activations, per-channel symmetric quantization for weights, and per-head asymmetric quantization for the KV cache. We fuse the projection matrices ${{\bm{U}}_{A}},{{\bm{U}}_{B}},{{\bm{U}}_{D}}$ into weights and apply GPTQ (Frantar et al., 2022) for weight quantization. To efficiently implement on-the-fly projections, ${{\bm{U}}_{D}}$ is a Hadamard matrix and ${\bm{U}}_{C}$ and its activations are quantized to 8-bit. The entire process, including obtaining projections and quantization, runs on a single NVIDIA A100 GPU; for Meta-Llama-3-8B, it takes 35 minutes. Additional details are in Appendix C.
Table 3: MMMU accuracy (higher is better) of vision language models when quantized using various approaches. For 4-bit data structures, our work ResQ and QUIK (Ashkboos et al., 2024b) keep $\nicefrac{{1}}{{8}}$ of channels in 8-bit.
Model
W/A/KV (bit) Method Qwen2-VL -2B-Instruct Qwen2-VL -7B-Instruct
16/16/16 Baseline 39.6 51.6
[…]
GPTQ 27.7 24.9
QUIK 26.3 28.9
QuaRot 24.0 24.5
4/4/4 ResQ 29.7 47.0
RTN 24.9 25.2
GPTQ 23.4 24.3
QUIK 28.4 26.4
QuaRot 26.5 24.5
4/8/4 ResQ 34.0 48.8
Table 4: Wikitext perplexity comparison of ResQ and baseline which keeps channels with high $l_{\infty}$-norm in high precision and uses rotation to reduce quantization error within high precision and low precision groups.
W/A/KV (bit) Method
Meta-Llama
-3-8B
Llama
-3.2-3B
@@ 5.2 Main Results
Language modeling, understanding, and reasoning tasks  We evaluate ResQ on tasks that test language modelling ability (perplexity on Wikitext), common sense reasoning ability (average 0-shot accuracy on the eight tasks listed in section 5.1) and language understanding (average 0-shot accuracy on MMLU). The results are presented in Table 1. We see that ResQ reduces the gap to 16-bit performance and outperforms the quantization baselines across all tasks on all models. Particularly, on Llama 3/3.2 family of models, ResQ outperforms SpinQuant by achieving 4-33% lower Wikitext perplexity, 0.1-5.4% better average 0-shot accuracy and a 1-14.5% better accuracy on MMLU benchmark without any additional training. For the Qwen-2.5 model family, all other baselines fail to achieve competitive results, and ResQ significantly outperforms them. Compared with QUIK, another mixed precision quantization approach, ResQ achieves 42-50% better Wikitext perplexity, 5.8-12.3% better average zero shot accuracy and 4.3-24.5% better MMLU accuracy over all models. Complete results on Llama and Qwen2.5 family of models are provided in Appendix D. Additionally, we provide comparison between baselines at W/A/KV = 4/8/4 bits for Llama families in Appendix E. Among all the set of results, ResQ maintains superior performance.
Generative tasks  We also test ResQ on tasks that require auto-regressive token generation including the GSM8K mathematical understanding benchmark, dialogue summarization benchmarks (qmsum and samsum) and code completion benchmark (repobench-p, Table 2). The goal of choosing these tasks is to evaluate the generation ability on a wide variety of domains. On the challenging GSM8K benchmark where QUIK fails to produce meaningful results, ResQ outperforms SpinQuant by 3.8% and 5.5% on the 8B and 3B parameter model respectively, closing the gap to the 16-bit baseline. On LongBench evaluation tasks, ResQ demonstrates competitive performance and outperforms SpinQuant without any additional training.
Multi-modal understanding  We benchmark the quantization approaches on vision language models (VLMs) by quantizing Qwen2 VL family and evaluating their performance on MMMU (Table 3, Yue et al. 2024). Only the language model is quantized while the vision encoder remains in 16-bit as the language model has many more parameters (over 10×10$\times$10 × for Qwen2-VL-7B-Instruct). ResQ outperforms baselines on both 2B and 7B models, achieving superior accuracy and demonstrating its generalizability. Results for individual MMMU tasks are provided in Appendix F.
Comparison against outliers with rotation baseline  A stronger baseline can be created combining existing quantization approaches. Like QUIK, one can find channels which consistently contain outliers and keep them in 8-bit while keep the remaining channels in low precision. And, the quantization of high/low precision groups can be improved using random rotations introduced in QuaRot. Compared with such a baseline which keeps channels with high $l_{\infty}$-norm in 8-bit, ResQ’s unique approach involves keeping coefficients along bases with high eigenvalues in 8-bit. We see in Table 4 that ResQ consistently outperforms such a strong baseline across various precisions of W/A/KV highlighting ResQ’s PCA driven theoretically optimal approach of choosing high precision components.
Table 5: Impact of different projections in ResQ. Evaluated by removing components and observing Wikitext perplexity.
@@ 5.3 Hardware performance
We implement the mixed-precision quantization using CUDA 11.8 and PyTorch. We use CUTLASS (Thakkar et al., 2023) to perform INT4 and INT8 GEMM operations on TensorCore. On NVIDIA RTX 3090 GPU, representative of resource constrained setting, we achieve about 1.61×1.61$\times$1.61 × to 3.03×3.03$\times$3.03 × speedup with ResQ over 16-bit baseline for a single decoder block tested across several language models (Figure 5). We observe higher speedup on larger models and smaller sequence lengths. Compared with INT4 implementation, ResQ kernel is only $14\%$ slower on average across models and sequence lengths demonstrating minimal overhead associated with mixed precision computations and on-the-fly projections introduced in ResQ.
Figure 5: Speedup of ResQ and INT4 kernel on single decoder block on NVIDIA RTX 3090 over 16-bit floating point baseline for batch size of 1.
@@ 5.4 Ablation studies
Projection bases  We evaluate the impact of different projections employed in ResQ by removing them and evaluating performance in Table 5. We see that removing ${{\bm{U}}_{D}}$ or ${{\bm{U}}_{A}}$ has a catastrophic impact on perplexity highlighting their importance. ${{\bm{U}}_{B}}$ and ${{\bm{U}}_{C}}$ which aid in quantization of KV cache have less severe impact when removed independently. But removing both of them leads to a non trivial increase in perplexity (particularly for Meta-Llama-3-8B and Llama-3.2-3B which employ grouped query attention).
@@
Calibration dataset  We change number of Wikitext calibration samples used to obtain projections and evaluate performance in Figure 6(b). For Meta-Llama-3-8B, MMLU accuracy increases with increasing samples and saturates beyond 128 samples. For Llama-3.2-3B, the trend is unclear with 512 samples achieving best performance.
@@ 6 Conclusion
We introduce ResQ, a novel mixed-precision, accelerator-friendly PTQ technique toward 4-bit quantization of large language models. ResQ projects weight, activation, and KV cache tensors to subspaces spanned by principal components, quantizing a low-rank ($\nicefrac{{1}}{{8}}$ of hidden dimension) high-variance subspace to 8-bit and the rest to 4-bit. ResQ outperforms both uniform- and mixed-precision quantization methods. We demonstrate the effectiveness of ResQ across a variety of tasks—including language modeling, language understanding, common-sense reasoning, language generation and multi modal understanding—using the Llama and Qwen models. Compared to SpinQuant, the strongest baseline, ResQ achieves up to 33% lower perplexity on the WikiText dataset without requiring any additional training and offers up to 3.03×3.03$\times$3.03 × speedup over the 16-bit baseline.
Impact Statement
ResQ is a significant step forward towards efficiently serving LLMs in resource-constrained, on-device scenarios, potentially expanding the application space for these models. Although our approach aims to make LLMs more accessible and widely used, it does not address the potential risks of misuse for malicious purposes. To mitigate these risks, a strong commitment to user data protection, clear ethical guidelines, and transparency mechanisms is essential.
Acknowledgements